### v1

# Experimental Demonstration of a Decentralized Electromagnetic Formation Flying Control Using Alternating Magnetic Field Forces

Abstract
Electromagnetic formation flying (EMFF) is challenging due to the complex coupling between the electromagnetic fields generated by each satellite in the formation. To address this challenge, this article uses alternating magnetic field forces (AMFF) to decouple the electromagnetic forces between each pair of satellites. Each satellite’s electromagnetic actuation system is driven by a sum of amplitude-modulated sinusoids, where amplitudes are controlled to achieve desired forces between each pair of satellites. The main contribution of this article is a 3-satellite experimental demonstration of decentralized closed-loop EMFF using AMFF. To our knowledge, this is the first demonstration of AMFF with at least 3 satellites in open or closed loop. This is noteworthy because the coupling challenges of EMFF are only present with more than 2 satellites, and thus, a formation of at least 3 is necessary to evaluate the effectiveness of AMFF. The experiments are conducted on a ground-based testbed consisting of 3 electromagnetically actuated satellites on linear air tracks. The closed-loop experimental results are compared with behavior from numerical simulations.
1 Introduction
Electromagnetic formation flying (EMFF) is accomplished using electromagnetic coils onboard satellites in a formation. Each satellite’s coils generate a magnetic field, which interacts with the magnetic fields of the other satellites to create magnetic field forces. These magnetic field forces can control the relative positions of satellites [10, 12, 14]. An advantage of EMFF is that the power source is renewable, while conventional actuation systems such as propellant thrusters eventually deplete.
EMFF for a single pair of satellites where the coils are driven by direct current (DC) is addressed in [6, 11]. Experiments using DC-driven coils are presented in [11]; this includes open-loop experiments as well as position-hold and trajectory-following experiments. In addition, feasibility experiments onboard the International Space Station are presented in [14], where attractive and repulsive forces were generated between 2 DC-driven coils.
EMFF for more than 2 satellites is challenging because the intersatellite forces are nonlinear functions of the magnetic moments generated by all satellites in the formation as well as the relative positions of all satellites. In other words, there is complex coupling between the electromagnetic fields generated by all satellites, and this coupling ultimately leads to intersatellite forces between every pair of satellites. EMFF control for more than 2 satellites is considered in [3, 15, 4]. However, these approaches require centralization of all measurement information and solving nonconvex constrained optimization problems that do not scale well to a large number of satellites.
A different approach to EMFF for more than 2 satellites is to use alternating magnetic field forces (AMFF) to address the challenge of coupling between the electromagnetic fields. The key idea of AMFF is that a pair of alternating (e.g., sinusoidal) magnetic moments results in a nonzero average interaction force between the pair of satellites if and only if those alternating magnetic moments have the same frequency [21, 13].
A decentralized EMFF method using AMFF is presented in [1]. In this method, each satellite has access to measurements of its position and velocity relative to only its local neighbor satellites. Then, each satellite uses a sum of frequency-multiplexed sinusoidal magnetic moments to achieve desired intersatellite forces between each pair of satellites. Other examples of EMFF using AMFF include [7, 22, 2, 18, 17, 19]. For example, [7, 22] uses AMFF for leader following. Relative position and attitude control with AMFF is addressed using a centralized algorithm in [19] and using a decentralized algorithm in [2]. In [18, 17], AMFF is combined with traditional reaction wheels in a centralized algorithm for relative position and attitude control.
Experimental evaluations of EMFF with AMFF are limited to [13, 2, 20]. Open-loop experiments demonstrating the AMFF approach are presented in [13]. Closed-loop single-degree-of-freedom experiments with a pair of satellites on linear air-tracks are presented in [2, 20]. Since the coupling challenges of EMFF are only present with more than 2 satellites, a formation of at least 3 satellites is necessary to evaluate the effectiveness of AMFF. However, to our knowledge, the existing literature does not include any experiments using AMFF with more than 2 satellites
The main contribution of this article is an experimental demonstration of closed-loop EMFF using AMFF with 3 satellites. To our knowledge, this is the first demonstration of AMFF with at least 3 satellites in open or closed loop. This work demonstrates EMFF with 3 satellites on a linear air track, where the coil of the center satellite is driven by a sum of 2 AC currents and each outer satellite is driven by one of the AC currents. As a result, intersatellite forces are generated between the center satellite and each outer satellite. This article adopts the piecewise-sinusoidal magnetic moment approach in [1] to address the challenges of coupling between the electromagnetic fields; this approach is reviewed in Section 4. Then, Section 5 presents the closed-loop decentralized EMFF algorithm, which is adopted from [1] but with a simplified approach for the control amplitude allocation. Section 6 describes the experimental testbed and modifies the EMFF control to account for hardware limitations such as sensor noise and amplitude saturation in the control current. Finally, experimental results are presented in Section 7. First, we present open- and closed-loop experiments with 2 satellites to demonstrate the feasibility of EMFF with AMFF. Finally, we present closed-loop experiments with 3 satellites, demonstrating frequency-multiplexed AMFF.
2 Notation
Physical vectors are denoted with bold symbols, for example, $\mathbf{r}$. The magnitude of $\mathbf{r}$ is denoted by $|\mathbf{r}|$. A frame is a collection of mutually-orthogonal physical unit vectors.
Let $\mathbb{N}$ denote the set of nonnegative integers. We let $\mathcal{I}\triangleq\{1,...,n\}$, where $n\in\mathbb{N}$ is the number of satellites in the formation, and $\mathcal{P}\triangleq\{(i,j)\in\mathcal{I}\times\mathcal{I}:i\neq j\}$, which is the set of ordered pairs. Unless otherwise stated, all statements in this paper that involve the subscript $i$, ijij, and $k$ are for all $i\in\mathcal{I}$, all $(i,j)\in\mathcal{P}$, and all $k\in\mathbb{N}$.
Figure 1: Each satellite is equipped with an electromagnetic actuation system consisting of three orthogonal coils. The relative positions of the satellites are controlled by the interaction of equal-and-opposite intersatellite forces produced by the actuation systems.
3 Dynamics
Consider a system of $n$ satellites, where each satellite has mass $m$. The position $\mathbf{r}_{i}$ locates the mass center of satellite $i$ relative to the origin of an inertial frame $\mathcal{F}$ which consists of the right-handed set of mutually orthogonal unit vectors $\begin{bmatrix}\mathbf{i}&\mathbf{j}&\mathbf{k}\end{bmatrix}$. The velocity $\mathbf{v}_{i}$ and acceleration $\dot{\mathbf{v}}_{i}$ are the first and the second time-derivatives of $\mathbf{r}_{i}$ with respect to $\mathcal{F}$. The relative position $\mathbf{r}_{ij}\triangleq\mathbf{r}_{i}-\mathbf{r}_{j}$ locates the mass center of satellite $i$ relative to the mass center of satellite $j$. The relative velocity $\mathbf{v}_{ij}$ is the time derivative of $\mathbf{r}_{ij}$ with respect to $\mathcal{F}$. The satellite system is shown in Figure 1.
Each satellite has an electromagnetic actuation system (i.e., multiple electromagnetic coils) that creates a magnetic field. These magnetic fields interact to produce intersatellite forces which control the satellites’ relative positions. The coils are modeled as magnetic dipoles, and the resulting intersatellite force applied to satellite $i$ by satellite $j$ is
$\mathbf{F}_{ij}=\frac{3\mu_{0}}{4\pi|\mathbf{r}_{ij}|^{4}}\mathbf{f}(\mathbf{r}_{ij},\mathbf{u}_{i},\mathbf{u}_{j}),$ (1)
where $\mu_{0}$ is the vacuum permeability constant, $\mathbf{u}_{i}$ is the magnetic moment of satellite $i$, and
$\mathbf{f}(\mathbf{r}_{ij},\mathbf{u}_{i},\mathbf{u}_{j}$ $)\triangleq\mathopen{}\mathclose{{\left(\mathbf{u}_{j}\cdot\frac{\mathbf{r}_{ij}}{|\mathbf{r}_{ij}|}}}\right)\mathbf{u}_{i}+\mathopen{}\mathclose{{\left(\mathbf{u}_{i}\cdot\frac{\mathbf{r}_{ij}}{|\mathbf{r}_{ij}|}}}\right)\mathbf{u}_{j}$
$+\mathopen{}\mathclose{{\left[(\mathbf{u}_{i}\cdot\mathbf{u}_{j})-5\mathopen{}\mathclose{{\left(\mathbf{u}_{i}\cdot\frac{\mathbf{r}_{ij}}{|\mathbf{r}_{ij}|}}}\right)\mathopen{}\mathclose{{\left(\mathbf{u}_{j}\cdot\frac{\mathbf{r}_{ij}}{|\mathbf{r}_{ij}|}}}\right)}}\right]\frac{\mathbf{r}_{ij}}{|\mathbf{r}_{ij}|}.$ (2)
The magnetic moment $\mathbf{u}_{i}$, which is the control of satellite $i$, is a function of the current supplied to the electromagnetic coils. Since $\mathbf{r}_{ij}=-\mathbf{r}_{ji}$, it follows from (1) and (2) that $\mathbf{F}_{ij}=-\mathbf{F}_{ji}$, that is, the force applied to satellite $j$ by $i$ is equal and opposite the force applied to $i$ by $j$. See [3] and [15] for more information on the force model (1) and (2).
The translational dynamics of satellite $i$ are
$\dot{\mathbf{v}}_{i}\triangleq\frac{c_{0}}{m}\sum_{j\in\mathcal{I}\setminus\{i\}}\frac{1}{|\mathbf{r}_{ij}|^{4}}\mathbf{f}(\mathbf{r}_{ij},\mathbf{u}_{i},\mathbf{u}_{j}),$ (3)
where $c_{0}\triangleq 3\mu_{0}/(4\pi)$. Since $\mathbf{F}_{ij}=-\mathbf{F}_{ji}$, it follows from (3) that $\sum_{i\in\mathcal{I}}\dot{\mathbf{v}}_{i}(t)=0$, which implies that the linear momentum of the system of satellites is conserved. Thus, the intersatellite electromagnetic forces can be used to alter relative positions, but they have no effect on the overall mass center of the satellites.
Let the constant $\mathbf{d}_{ij}$ be the desired position of satellite $i$ relative to satellite $j$, where $\mathbf{d}_{ij}=-\mathbf{d}_{ji}$. The objective is to design controls that rely on feedback of neighboring satellites and yield $\lim_{t\to\infty}\mathbf{r}_{ij}(t)=\mathbf{d}_{ij}$.
4 Time-Averaged Dynamics with Piecewise-
Sinusoidal Controls
This section reviews the piecewise-sinusoidal magnetic-moment approach in [1], which is used to address the coupling that occurs between electromagnetic fields. In this approach, each satellite uses a piecewise-sinusoidal magnetic moment $\mathbf{u}_{i}$ that is the sum of n1n-1 sinusoids with n1n-1 unique frequencies. Each unique frequency is common to only one pair of satellites. Thus, there are a total of n(n1)/2n(n-1)/2 unique frequencies, and the amplitudes of each common-frequency pair of sinusoids are selected to prescribe the average intersatellite force between the associated satellite pair.
Let $\omega_{ij}\geq 0$ be the interaction frequency, where $\omega_{ij}=\omega_{ji}$ is unique. Next, let $T>0$ be a common multiple of $\{2\pi/\omega_{ij}:(i,j)\in{\mathcal{P}}\}$. For all $k\in\mathbb{N}$ and $t\in[kT,kT+T)$, we consider the piecewise-sinusoidal control
$\mathbf{u}_{i}(t)=\sum_{j\in\mathcal{I}\setminus\{i\}}\mathbf{p}_{ij,k}\sin\omega_{ij}t$ (4)
where the amplitudes $\{\mathbf{p}_{ij,k}\}^{\infty}_{k=0}$ are the control variables.
The following result characterizes the decoupling feature of the piecewise-sinusoidal control (4). This result is from [1, Proposition 1] and is a consequence of averaging the product of sinusoids over an integer number of periods.
Proposition 1.
Let $i,j\in\mathbb{N}$, and let $\mathbf{r}$ be a constant. Then, for all $k\in\mathbb{N}$,
$\frac{1}{T}\int^{kT+T}_{kT}\mathbf{f}(\mathbf{r},\mathbf{u}_{i}(t),\mathbf{u}_{j}(t))\ \text{d}t\triangleq\frac{1}{2}\mathbf{f}(\mathbf{r},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k}).$
We use Proposition 1 to derive a time-averaged model of the dynamics (3). Integrating (3) over the interval $[kT,kT+T]$ yields
[…]
where the average intersatellite force is
$\bar{\mathbf{F}}_{ij}(k)\triangleq\frac{1}{T}\int_{kT}^{kT+T}\frac{c_{0}}{|\mathbf{r}_{ij}(t)|^{4}}\mathbf{f}(\mathbf{r}_{ij}(t),\mathbf{u}_{i}(t),\mathbf{u}_{j}(t))\ \text{d}t.$
We define the approximate average intersatellite force
$\hat{\mathbf{F}}_{ij,k}\triangleq\frac{1}{T}\int_{kT}^{kT+T}\frac{c_{0}}{|\mathbf{r}_{ij,k}|^{4}}\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{u}_{i}(t),\mathbf{u}_{j}(t))\ {\rm d}t,$ (6)
which is an approximation of $\bar{\mathbf{F}}_{ij,k}$ obtained by replacing $\mathbf{r}_{ij}(t)$ with its sampling $\mathbf{r}_{ij,k}\triangleq\mathbf{r}_{ij}(kT)$. Propositon 1 implies that
$\hat{\mathbf{F}}_{ij,k}=\frac{c_{0}}{2|\mathbf{r}_{ij,k}|^{4}}\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k}).$ (7)
Note that $(\hat{\mathbf{F}}_{ij,k},\hat{\mathbf{F}}_{ji,k})$ depends on the amplitude control pair $(\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})$, which does not affect the approximate average intersatellite force between any other pair of satellites.
5 EMFF Control Algorithm
5.1 Desired Intersatellite Forces
The intersatellite feedback structure is described using an undirected graph. The vertex set of the undirected graph is $\mathcal{I}$. Let $\mathcal{E}\subset\mathcal{I}\times\mathcal{I}$ be the edge set. We assume that the undirected graph $\mathcal{G}=(\mathcal{I},\mathcal{E})$ is connected. The neighbor set of satellite $i$ is $\mathcal{N}_{i}\triangleq\{j\in\mathcal{I}:(i,j)\in\mathcal{E}\}$, and satellite $i$ has access to $\{\mathbf{r}_{ij,k}\}_{j\in\mathcal{N}_{i}}$ and $\{\mathbf{v}_{ij,k}\}_{j\in\mathcal{N}_{i}}$ for feedback, where $\mathbf{v}_{ij,k}\triangleq\mathbf{v}_{i}(kT)-\mathbf{v}_{j}(kT)$.
Define the desired value for the intersatellite force function
$\mathbf{f}^{*}_{ij,k}\triangleq-\frac{2m|\mathbf{r}_{ij,k}|^{4}}{c_{0}}\alpha_{ij}((\mathbf{r}_{ij,k}-\mathbf{d}_{ij})+\beta\mathbf{v}_{ij,k})$ (8)
where $\alpha_{ij}=\alpha_{ji}\geq 0$ and $\beta>0$ are feedback gains. If $(i,j)\notin\mathcal{E}$, then $\alpha_{ij}=0$; otherwise, $\alpha_{ij}>0$.
If $\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})=\mathbf{f}^{*}_{ij,k}$, then it follows from 7 the approximate average intersatellite force is
$\hat{\mathbf{F}}_{ij,k}=-m\alpha_{ij}\mathopen{}\mathclose{{\left[(\mathbf{r}_{ij,k}-\mathbf{d}_{ij})+\beta\mathbf{v}_{ij,k}}}\right].$ (9)
In this case, [1, Theorem 5] shows how to select the gains $\alpha_{ij}$ and $\beta$ such that the formation error $\mathbf{r}_{ij,k}-\mathbf{d}_{ij}$ tends to zero for $\hat{\mathbf{F}}_{ij,k}=\hat{\mathbf{F}}_{ij,k}$. The approximate average intersatellite force 9 is based on linear consensus for sampled-data double integrators [5]. This method uses feedback to determine approximate average intersatellite forces $\hat{\mathbf{F}}_{ij,k}$ that create virtual linear springs and dashpots between satellite $i$ and $j$.
5.2 Control Amplitude Pair To Achieve Prescribed Intersatellite Force
This section provides a construction for the amplitude pair $(\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})$ such that $\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})=\mathbf{f}^{*}_{ij,k}$. This construction is presented in [9, 8] for the system resolved in an inertial frame. Here, we present the construction coordinate free.
Define
$g_{\mathrm{r}}(\mathbf{r},\mathbf{f}^{*})$ $\triangleq-\frac{\operatorname{sgn}(\mathbf{r}\cdot\mathbf{f}^{*})}{2}\mathopen{}\mathclose{{\left(\frac{|\mathbf{r}\cdot\mathbf{f}^{*}|+\Phi_{1}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}|}}}\right)^{\frac{1}{2}},$ (10)
$g_{\mathrm{rf}}(\mathbf{r},\mathbf{f}^{*})$ $\triangleq\frac{1}{\sqrt{2}}\mathopen{}\mathclose{{\left(\frac{-|\mathbf{r}\cdot\mathbf{f}^{*}|+\Phi_{2}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}|}}}\right)^{\frac{1}{2}},$ (11)
$h_{\mathrm{r}}(\mathbf{r},\mathbf{f}^{*})$ $\triangleq\frac{1}{2}\mathopen{}\mathclose{{\left(\frac{|\mathbf{r}\cdot\mathbf{f}^{*}|+\Phi_{2}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}|}}}\right)^{\frac{1}{2}},$ (12)
$h_{\mathrm{rf}}(\mathbf{r},\mathbf{f}^{*})$ $\triangleq-\frac{\operatorname{sgn}(\mathbf{r}\cdot\mathbf{f}^{*})}{\sqrt{2}}\mathopen{}\mathclose{{\left(\frac{-|\mathbf{r}\cdot\mathbf{f}^{*}|+\Phi_{1}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}|}}}\right)^{\frac{1}{2}},$ (13)
where
$\Phi_{1}(\mathbf{r},\mathbf{f}^{*})$ $\triangleq\sqrt{|\mathbf{r}\times\mathbf{f}^{*}|^{2}+|\mathbf{r}|^{2}|\mathbf{f}^{*}|^{2}},$ (14)
$\Phi_{2}(\mathbf{r},\mathbf{f}^{*})$ $\triangleq(2-\operatorname{sgn}(\mathbf{r}\cdot\mathbf{f}^{*})^{2})\Phi_{1}(\mathbf{r},\mathbf{f}^{*}).$ (15)
Next, define
$\mathbf{g}(\mathbf{r},\mathbf{f}^{*})$ $\triangleq\begin{cases}\frac{g_{\mathrm{r}}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}|}\mathbf{r}+\frac{g_{\mathrm{rf}}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}||\mathbf{r}\times\mathbf{f}^{*}|}((\mathbf{r}\times\mathbf{f}^{*})\times\mathbf{r}),&\mathbf{r}\times\mathbf{f}^{*}\neq\mathbf{0},\\ \frac{g_{\mathrm{r}}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}|}\mathbf{r},&\mathbf{r}\times\mathbf{f}^{*}=\mathbf{0},\end{cases}$ (16)
$\mathbf{h}(\mathbf{r},\mathbf{f}^{*})$ $\triangleq\begin{cases}\frac{h_{\mathrm{r}}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}|}\mathbf{r}+\frac{h_{\mathrm{rf}}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}||\mathbf{r}\times\mathbf{f}^{*}|}((\mathbf{r}\times\mathbf{f}^{*})\times\mathbf{r}),&\mathbf{r}\times\mathbf{f}^{*}\neq\mathbf{0},\\ \frac{h_{\mathrm{r}}(\mathbf{r},\mathbf{f}^{*})}{|\mathbf{r}|}\mathbf{r},&\mathbf{r}\times\mathbf{f}^{*}=\mathbf{0}.\end{cases}$ (17)
The following result shows that 16 and 17 is a construction for the amplitude pair such that $\mathbf{f}$ takes on a prescribed value $\mathbf{f}^{*}$. The proof is in A.
Proposition 2.
For all $\mathbf{f}^{*}$ and all $\mathbf{r}\neq\mathbf{0}$,
$\mathbf{f}(\mathbf{r},\mathbf{g}(\mathbf{r},\mathbf{f}^{*}),\mathbf{h}(\mathbf{r},\mathbf{f}^{*}))=\mathbf{f}^{*}.$
Proposition 2 implies that we can use 16 and 17 to select the amplitude pair $(\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})$ such that the average intersatellite force equals the desired value 8. Specifically, we let the amplitudes be
[…]
and it follows from Proposition 2 that $\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})=\mathbf{f}^{*}_{ij,k}$. In other words, the amplitude pair $(\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})$ given by (18) achieves the desired intersatellite force $\mathbf{f}^{*}_{ij,k}$.
6 Experiment Setup
This section describes a one-dimensional (1D) experimental platform for testing the EMFF algorithm from Sections 4 and 5, specializes the algorithm to 1D, and presents modifications for implementation.
Figure 2: Experimental platform. Three EAS units sit on the linear air tracks, where the air is supplied by the air sources to create low-friction motion.
Figure 3: Front view, side view, and top view of the EAS unit on the air track, displaying the power electronics, the arduino, and the laser range-finding sensor.
6.1 Experimental Platform
The experimental platform is shown in Fig. 2. The testbed includes four Cylone Pro air blowers, two Eisco PH0362A Linear Air Tracks, and three custom-designed one-dimensional electromagnetic actuation system (EAS) units. Each air track is attached to one Cylone Pro air blower. The air blowers supply the required air pressure to the linear air tracks. The EAS units sit on four air track gliders to allow for low friction motion of the units on the air track. The air tracks are mounted on an aluminum jig-plate which is mounted on a flat and level surface table.
Each EAS unit has an electromagnetic coil and an electrical and electronics platform. Each electromagnetic coil has an area of $A=0.1\pi^{2}$ m2 and consists of $N=500$ turns of 22 gauge copper wire. The electrical and electronics platforms include all components needed to supply power to the electromagnetic coils. The power electronics consist of two lithium-ion polymer batteries, two Vicor DC-DC converters, and a Copley Controls Junus JSP-090-10 amplifier, as shown in Figure 3.
Each EAS unit has an Arduino Due microcontroller which is responsible for data acquisition and feedback control. An XBee radio frequency (RF) module is connected to an Arduino wireless secure digital (SD) shield, which is attached to the microcontroller. A coordinator Xbee RF module is also connected to a computer, enabling communication between the computer and the EAS units through a Digi XCTU application.
The outer EAS units have a single laser ranging module (STM VL53L0X) connected to the microcontroller, while the middle EAS unit has 2 laser-ranging modules on either side. The laser ranging modules are used to measure the unit’s position relative to the other unit. The laser ranging module is mounted on an aluminum plate, which forms the base of each EAS unit. The laser ranging module faces a white reflector that is mounted to the other EAS unit.
6.2 Control in 1D with Current Implementation
This section specializes to 1D, where all motion is in the $\mathbf{i}$ direction. Specifically, let $r_{ij,k}\in\mathbb{R}$, $v_{ij,k}\in\mathbb{R}$, and $d_{ij}\in\mathbb{R}$ be such that
$\mathbf{r}_{ij,k}=r_{ij,k}\mathbf{i},\quad\mathbf{v}_{ij,k}=v_{ij,k}\mathbf{i},\quad\mathbf{d}_{ij}=d_{ij}\mathbf{i},$ (19)
and it follows from 8 that
$\mathbf{f}^{*}_{ij,k}=f^{*}_{ij,k}\mathbf{i},$ (20)
where
$f^{*}_{ij,k}\triangleq-\frac{2m|r_{ij,k}|^{4}}{c_{0}}\alpha_{ij}((r_{ij,k}-d_{ij})+\beta v_{ij,k}).$ (21)
For implementation, we do not control magnetic moment 4 directly. Instead, the input is the current $I_{i}:\mathopen{}\mathclose{{\left[0,\infty}}\right)\rightarrow\mathbb{R}$ in the $i$th coil, which is
$I_{i}(t)=\sum_{{j}\in\mathcal{N}_{i}}I_{ij,k}\sin\omega_{ij}t,$ (22)
where
$I_{ij,k}\triangleq I^{*}(r_{ij,k},f^{*}_{ij,k},i-j),$ (23)
and
$I^{*}(r_{ij,k},f^{*}_{ij,k},i-j)\triangleq\begin{cases}-\dfrac{\text{sgn}(f^{*}_{ij,k})}{NA}\sqrt{\dfrac{|f^{*}_{ij,k}|}{2}}\mathbf{,}&i-j<0,\\ \dfrac{\text{sgn}(r_{ij,k})}{NA}\sqrt{\dfrac{|f^{*}_{ij,k}|}{2}}\mathbf{,}&i-j>0.\end{cases}$ (24)
The current control (22) generates a piecewise sinusoidal magnetic moment with frequency $\omega_{ij}$ and amplitude
$p_{ij,k}=NAI_{ij,k}.$ (25)
Thus, (22) generates the magnetic moment 4, where $\mathbf{p}_{ij,k}=p_{ij,k}\mathbf{i}$, which is equal to 18 where $\mathbf{f}^{*}_{ij,k}=f^{*}_{ij,k}\mathbf{i}$ and $\mathbf{r}_{ij,k}=r_{ij,k}\mathbf{i}$.
To verify that $\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ij,k})=\mathbf{f}^{*}_{ij,k}$, note that since $\mathbf{r}_{ij,k}=r_{ij,k}\mathbf{i}$, and $\mathbf{p}_{ij,k}=p_{ij,k}\mathbf{i}$, it follows from (2) that
$\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})=f(r_{ij,k},p_{ij,k},p_{ji,k})\mathbf{i},$ (26)
where
$f(r_{ij,k},p_{ij,k},p_{ji,k})\triangleq-2\operatorname{sgn}(r_{ij,k})p_{ij,k}p_{ji,k}.$ (27)
Substituting 23, 24, and 25 into 27 yields $f(r_{ij,k},p_{ij,k},p_{ji,k})=f^{*}_{ij,k}$, which combined with 20 and 26 demonstrates that $\mathbf{f}(\mathbf{r}_{ij,k},\mathbf{p}_{ij,k},\mathbf{p}_{ji,k})=\mathbf{f}^{*}_{ij,k}$.
6.3 Control Modification to Address Hardware Limitations
We introduce 2 modifications to the current control 21, 22, 23, and 24. First, we introduce integral action to the desired force 21 to address steady-state error. Second, we introduce modifications to the current allocation 23 to address current amplitude saturation. This section redefines $f_{ij,k}^{*}$ and $I_{ij,k}$ given by 21 and 23 based on these modifications.
Let $\epsilon_{0}>0$ and $\epsilon_{1}>\epsilon_{0}$, and let the integrator state $\xi_{ij,k}$ satisfy
$\xi_{ij,k}$ $\triangleq\begin{cases}\xi_{ij,k-1}+r_{ij,k}-d_{ij},&|r_{ij,k}-d_{ij}|\in(\epsilon_{0},\epsilon_{1}),\\ 0,&|r_{ij,k}-d_{ij}|\notin(\epsilon_{0},\epsilon_{1}),\end{cases}$ (28)
where $\xi_{ij,0}\in\mathbb{R}$ is the initial condition. Note that 28 is an integrator if the formation error $r_{ij,k}-d_{ij}$ has small magnitude (i.e., in the interval $(\epsilon_{0},\epsilon_{1})$). Otherwise, the integrator is disabled. Then, define the modified desired formation force
$f^{*}_{ij,k}\triangleq-\frac{2m|r_{ij,k}|^{4}}{c_{0}}\mathopen{}\mathclose{{\left(\alpha_{ij}\mathopen{}\mathclose{{\left(\mathopen{}\mathclose{{\left(r_{ij,k}-d_{ij}}}\right)+\beta v_{ij,k}}}\right)+\rho_{ij}\xi_{ij,k}}}\right).$ (29)
where $\rho_{ij}=\rho_{ji}\geq 0$.
Since the intersatellite force depends on the product of $I^{*}(r_{ij,k},f^{*}_{ij,k},i-j)$ and $I^{*}(r_{ji,k},f^{*}_{ji,k},j-i)$, equal force can be achieved if one amplitude is increased and the other is decreased proportionally. Thus, we allocate more control authority to satellites with fewer neighbors. To do so, we multiply $I^{*}(r_{ij,k},f^{*}_{ij,k},i-j)$ with $\gamma_{ij}$ where $\gamma_{ij}>0$ and $\gamma_{ji}=1/\gamma_{ij}$.
Next, define
$\bar{I}_{i,k}\triangleq\max_{t\in[kT,kT+T)}\mathopen{}\mathclose{{\left|\sum_{j\in\mathcal{N}_{i}}\gamma_{ij}I^{*}(r_{ij,k},f^{*}_{ij,k},i-j)\sin\omega_{ij}t}}\right|,$
which is the maximum amplitude of the sum of sinusoidal currents over $t\in[kT,kT+T)$. Since the current is limited by the power capability of the EAS units, we modify $I_{ij,k}$ such that $\bar{I}_{i,k}$ does not exceed the maximum allowable current $\bar{I}>0$. The current implemented on the $i$th coil is 22, where the current amplitude is
$I_{ij,k}\triangleq\begin{cases}\gamma_{ij}I^{*}(r_{ij,k},f^{*}_{ij,k},i-j),&\quad\bar{I}_{i,k}\leq\bar{I},\\ \gamma_{ij}\dfrac{\bar{I}}{\bar{I}_{i,k}}I^{*}(r_{ij,k},f^{*}_{ij,k},i-j),&\quad\bar{I}_{i,k}>\bar{I}.\end{cases}$ (30)
Then, it follows from 22 and 30 that
$\max_{t\in[kT,kT+T)}|I_{i}(t)|\leq\bar{I},$ (31)
which implies that the current implemented is limited to the maximum allowable current.
Next, we show that the resultant prescribed force is a scaling of the unconstrained prescribed force $f^{*}_{ij,k}$. Specifically, substituting 24, 25, and 30 into 27 yields
$f(r_{ij,k},p_{ij,k},p_{ji,k})=\frac{\bar{I}^{2}}{(\max\{\bar{I},\bar{I}_{i,k}\})(\max\{\bar{I},\bar{I}_{j,k}\})}f^{*}_{ij,k}.$
Furthermore, 6 and 7 imply that the approximate average intersatellite force generated is equal to $\hat{\mathbf{F}}_{ij,k}=\hat{F}(r_{ij,k},p_{ij,k},p_{ji,k})\mathbf{i}$, where
$\hat{F}(r_{ij,k},p_{ij,k},p_{ji,k})\triangleq\frac{c_{0}}{2|r_{ij,k}|^{4}}f(r_{ij,k},p_{ij,k},p_{ji,k}).$ (32)
6.4 Estimation of Relative Position and Relative Velocity
Let $q_{ij,k}$ be the measurement of the relative position at time $kT$, which is taken on satellite $i$ using the onboard sensor. For each satellite, we design Kalman filters, where each Kalman filter uses $q_{ij,k}$ to obtain the estimates $\hat{r}_{ij,k}$ and $\hat{v}_{ij,k}$ of $r_{ij,k}$ and $v_{ij,k}$ for $j\in\mathcal{N}_{i}$.
It follows from 5, 6, and 7 that $\hat{r}_{ij,k}$ approximately satisfies sampled-data double integrator dynamics where the input is
$\nu_{ij,k}\triangleq$ $\frac{1}{m}\mathopen{}\mathclose{{\left(\sum_{g\in\mathcal{N}_{i}}\hat{F}(\hat{r}_{ig,k},p_{ig,k},p_{gi,k})}}\right$.
$-\mathopen{}\mathclose{{\left.\sum_{h\in\mathcal{N}_{j}}\hat{F}(\hat{r}_{jh,k},p_{jh,k},p_{hj,k})}}\right).$ (33)
Since satellite $i$ does not have information of all forces acting on satellite $j$, we consider only the intersatellite forces applied to satellite $j$ by the common neighbors of satellites $i$ and $j$. Thus, the estimate of $\nu_{ij,k}$ computed onboard satellite $i$ is
$\hat{\nu}_{ij,k}\triangleq$ $\frac{1}{m}\mathopen{}\mathclose{{\left(\sum_{g\in\mathcal{N}_{i}}\hat{F}(\hat{r}_{ig,k},p_{ig,k},p_{gi,k})}}\right$.
$-\mathopen{}\mathclose{{\left.\sum_{h\in\{i\}\cup(\mathcal{N}_{i}\cap\mathcal{N}_{j})}\hat{F}(\hat{r}_{jh,k},p_{jh,k},p_{hj,k})}}\right).$ (34)
Let $V>0$ be the variance of the noise associated with the measurement $q_{ij,k}$. Then for all $k$, the filtered relative position $\hat{r}_{ij,k}$ and relative velocity $\hat{v}_{ij,k}$ are given by
$\begin{bmatrix}\hat{r}_{ij,k}\\ \hat{v}_{ij,k}\end{bmatrix}$= $(A-LCA)\begin{bmatrix}\hat{r}_{ij,k-1}\\ \hat{v}_{ij,k-1}\end{bmatrix}+(B-LCB)\hat{\nu}_{ij,k-1}$
$+Lq_{ij,k},$ (35)
where
$A$ $\triangleq\begin{bmatrix}1&T\\ 0&1\end{bmatrix},\quad B\triangleq\begin{bmatrix}0.5T^{2}\\ T\end{bmatrix},\quad C\triangleq\begin{bmatrix}1&0\end{bmatrix}$, (36)
and
$L\triangleq PC^{\text{T}}(CPC^{\text{T}}+V)^{-1},$ (37)
where $P$ is the solution to the discrete algeabric Ricatti equation
$A^{\text{T}}PA-P-A^{\text{T}}PC(C^{\text{T}}PC+V)^{-1}C^{\text{T}}PA+W=0,$ (38)
where $W\triangleq wBB^{\mathrm{T}}$ and $w>0$ is the variance of the disturbance force.
7 Experimental Results and Discussion
We present results from open-loop and closed-loop experiments using the EAS units described in Section 6.1 with the control algorithm presented in Section 6.2 with the modifications presented in Section 6.3. The plots of the experiments include $|q_{ij,k}|$, $|\hat{r}_{ij,k}|$, $\hat{v}_{ij,k}$, $I_{ij,k}$, and $\hat{F}(r_{ij,k},p_{ij,k},p_{ji,k})$. The Kalman filter is initialized with $\hat{r}_{ij,0}=q_{ij,0}$ and $\hat{v}_{ij,0}=0$ m/s.
7.1 2 Satellite Experiments
We present 2 open-loop experiments and 3 closed-loop experiments with 2 satellites. The left satellite is numbered 1 and the right satellite is numbered 2. The actuation frequency are $\omega_{12}=\omega_{21}=40\pi$ rad/s, and the control is implemented at $T=0.1$ s. Using experimental data, we estimate the variances $V=1.2\times 10^{-6}$ $\mathrm{m}^{2}$ and $w=5\times 10^{-6}$ $\mathrm{m}^{2}/\mathrm{s}^{4}$. It follows from 36, 37, and 38 that $L=\begin{bmatrix}0.0942&0.0466\end{bmatrix}^{\mathrm{T}}.$
First, we present open-loop experiments to demonstrate that the experimental testbed is capable of generating attractive and repulsive forces.
Experiment 1.
This experiment demonstrates open-loop attraction. Figure 4 shows the experimental results, where the first row shows the raw relative-position measurements $|q_{ij}|$, the second and third rows show the filtered measurements $|\hat{r}_{ij}|$ and $\hat{v}_{ij}$, the fourth row shows the current $I_{ij}$, and the fifth row shows the average intersatellite force $\hat{F}(r_{ij,k},p_{ij,k},p_{ji,k})$. The first column shows the data measured and computed onboard satellite 1, and the second column shows the data measured and computed onboard satellite 2.
The control is turned on at $t=5$ s with amplitudes $I_{12}=1$ A and $I_{21}=1$ A (sinusoidal currents are in phase). Then, $|\hat{r}_{ij}|$ decreases from 0.508 m to 0.38 m over the next 15 s, which demonstrates attraction. There is a slight offset between the initial estimates $|\hat{r}_{12,0}|$ and $|\hat{r}_{21,0}|$, which resolves at smaller distance. This can be explained by sensor and reflector misalignment, and laser-ranging sensor bias at larger distances. $\triangle$
Figure 4: Open-loop attraction between two satellites where $I_{12,k}=1$ A and $I_{21,k}=1$ A for $t>5$ seconds. The sinusoidal currents on the satellites are in phase, which results in an attractive force.
Experiment 2.
This experiment demonstrates open-loop repulsion. Figure 5 shows that the control is turned on at $t=5$ s with amplitudes $I_{12}=-1$ A and $I_{21}=1$ A (sinusoidal currents are out of phase). Then, $|\hat{r}_{ij}|$ increases from 0.404 m to 0.56 m over the next 15 s, which demonstrates repulsion. $\triangle$
Figure 5: Open-loop repulsion between two satellites where $I_{12,k}=-1$ A and $I_{21,k}=1$ A for $t>5$ seconds. The sinusoidal currents on the satellites are out of phase, which results in a repulsive force.
Next, we present 3 experiments demonstrating closed-loop control. The control $(I_{12,k},I_{21,k})$ is implemented with $\alpha_{ij}=0.0158$, $\beta=6.89$, $\rho_{ij}=0$, and $\bar{I}=2.35$ A. Note that $\alpha_{ij}$ and $\beta$ are selected such that $\alpha_{ij}m/(c_{0}N^{2}A^{2})=812$ and $\alpha_{ij}\beta m/(c_{0}N^{2}A^{2})=5600$.
Experiment 3.
This experiment demonstrates closed-loop repulsion. The satellites start at a relative distance of approximately 0.4 m. The objective is to reach a desired relative distance $|d_{ij}|=0.45$ m. Figure 6 shows that $|\hat{r}_{ij}|$ settles to $|d_{ij}|=0.45$ m after approximately 20 s, which demonstrates that the satellites achieve formation. There is a sharp decrease in the measurement $|q_{12}|$ at $t=2.5$ s, which is due to a sensor dropout.
Next, we compare the experimental results with simulations of the closed-loop 3 and 4 where $\mathbf{p}_{ij,k}=p_{ij,k}\mathbf{i}$. We initialize the simulation using the same initial conditions as in the experiment. The simulated relative position includes zero-mean Gaussian-white sensor noise with variance $V=1.2\times 10^{-6}$ m2, which is the sensor noise variance estimated from experimental data. Then, the filtered relative position and velocity are obtained from the Kalman filter in Section 6.4, which is the method used in the experiment. Figure 7 shows the simulation results, which qualitatively match the experimental results in Figure 6.
To compare the experiment and simulation quantitatively, we use 4 metrics: overshoot $\hat{r}_{\mathrm{os}}$; settling time $T_{\mathrm{s}}$ (i.e. the time it takes $\hat{r}_{ij}$ to reach and stay within 1% of its final value); maximum value of $|\hat{F}(r_{ij,k},p_{ij,k},p_{ji,k})|$; and
$P\triangleq\sqrt{\frac{1}{N_{\mathrm{t}}}\sum_{k=0}^{N_{\mathrm{t}}}\mathopen{}\mathclose{{\left(\hat{F}(r_{ij,k},p_{ij,k},p_{ji,k})}}\right)^{2}},$ (39)
[…]
Figure 6: Experiment demonstrating closed-loop repulsion with 2 satellites, where the satellites start at approximately 0.4 m apart, and the desired relative position is $|d_{ij}|=0.45$ m.
Figure 7: Simulation of closed-loop repulsion with 2 satellites, where the satellites start at approximately 0.4 m apart, and the desired relative position is $|d_{ij}|=0.45$ m.
Figure 8: Experiment demonstrating closed-loop attraction with 2 satellites, where the satellites start at approximately 0.5 m apart, and the desired relative position is $|d_{ij}|=0.45$ m.
Figure 9: Simulation of closed-loop attraction with 2 satellites, where the satellites start at approximately 0.5 m apart, and the desired relative position is $|d_{ij}|=0.45$ m.
Table 1: Metrics for 3 and Simulation
3 Simulation
$(i,j)=(1,2)$ $(i,j)=(2,1)$ $(i,j)=(1,2)$ $(i,j)=(2,1)$
$\hat{r}_{\mathrm{os}}$ (cm) 0.76 0.95 0.55 0.47
$T_{\mathrm{s}}$ 21 21.1 18.3 18.3
[…]
$P$ $02\times 10^{-4} 2.2$$27\times 10^{-4}$ $1.43\times 10^{-4}$ $1.41\times 10^{-4}$
Experiment 4.
This experiment demonstrates closed-loop attraction. The satellites start at a relative distance of approximately $0.5$ m. The objective is to reach a desired relative distance $|d_{ij}|=0.45$ m. Figure 8 shows that $|\hat{r}_{ij}|$ settles to the desired relative position $|d_{ij}|=0.45$ m after approximately 20 s, which demonstrates that the satellites achieve formation. Note that the satellites move before the control is turned on at $t=5$ s. This can be explained by a slight impulse introduced at the start of the experiment and airflow from the airtracks creating disturbances.
[…]
Table 2: Metrics for 4 and Simulation
4 Simulation
$(i,j)=(1,2)$ $(i,j)=(2,1)$ $(i,j)=(1,2)$ $(i,j)=(2,1)$
$\hat{r}_{\mathrm{os}}$ (cm) 1.02 1.22 0.59 0.58
$T_{\mathrm{s}}$ 20.9 21 18.8 18.7
[…]
$P$ $2.38\times 10^{-4}$ $2.62\times 10^{-4}$ $1.66\times 10^{-4}$ $1.64\times 10^{-4}$
Figure 10: Experiment demonstrating multiple maneuvers with 2 satellites, where the satellites start at approximately 0.5 m apart. The initial desired relative position is $|d_{ij}|=0.4$ m, and at 60 s the desired relative position is switched to $|d_{ij}|=0.5$ m.
Figure 11: Simulation demonstrating multiple maneuvers with 2 satellites, where the satellites start at approximately 0.5 m apart. The initial desired relative position is $|d_{ij}|=0.4$ m, and at 60 s the desired relative position is switched to $|d_{ij}|=0.5$ m.
Experiment 5.
This experiment demonstrates multiple maneuvers in succession. The satellites start at a relative distance of approximately 0.50 m. The initial objective is to reach a desired relative distance $|d_{ij}|=0.4$ m. Next, the desired relative position is switched to $0.5$ m at $t=60$ s. Figure 10 shows that $|\hat{r}_{ij}|$ settles to $|d_{ij}|=0.40$ m at approximately 35 s, and $|\hat{r}_{ij}|$ settles to $|d_{ij}|=0.50$ m by approximately 90 s, which demonstrates that the satellites can achieve multiple formation maneuvers. Figure 11 shows the simulation results, which qualitatively matches the experimental results in Figure 10. $\triangle$
7.2 3 Satellite Closed-Loop Experiments
Next, we present 3 closed-loop experiments with $n=3$ satellites. The middle, left, and right satellites are numbered 1, 2, and 3. The neighbor sets are $\mathcal{N}_{1}=\{2,3\}$, $\mathcal{N}_{2}=\{1\}$, and $\mathcal{N}_{3}=\{1\}$. The interaction frequencies are $\omega_{12}=\omega_{21}=20\pi$ rad/s and $\omega_{13}=\omega_{31}=40\pi$ rad/s. We implement the currents $(I_{ij,k},I_{ji,k})$ given by (29)–(30) with $\epsilon_{0}=0.015$ m, $\epsilon_{1}=0.021$ m, $\alpha_{ij}=0.0158$, $\beta=7.38$, $\rho_{ij}=0.00136$, $\gamma_{12}=\gamma_{13}=0.8$ and $\gamma_{21}=\gamma_{31}=1.25$, and $\bar{I}=2.35$ A. Note that we select $\alpha_{ij}$, $\beta$, and $\rho_{ij}$ such that $\alpha_{ij}m/(c_{0}N^{2}A^{2})=812$, $\alpha_{ij}\beta m/(c_{0}N^{2}A^{2})=6000$, and $\rho_{ij}m/(c_{0}N^{2}A^{2})=70$.
Using experimental data, we estimate the noise variances $V=2\times 10^{-6}$ $\mathrm{m}^{2}$ and $w=5\times 10^{-6}$ $\mathrm{m}^{2}/\mathrm{s}^{4}$. It follows from 36, 37, and 38 that $L=\begin{bmatrix}0.1064&0.0598\end{bmatrix}^{\mathrm{T}}$.
Figure 12: Experiment demonstrating closed-loop repulsion between satellite 1 and 2 and closed-loop repulsion between satellite 1 and 3. Initially, satellites 1 and 2 are approximately 0.346 m apart, and satellites 1 and 3 are approximately 0.377 m apart. The desired relative positions are $d_{12}=0.42$ m, and $d_{13}=-0.45$ m.
Figure 13: Simulation demonstrating closed-loop repulsion between satellite 1 and 2 and closed-loop repulsion between satellite 1 and 3. Initially, satellites 1 and 2 are approximately 0.346 m apart, and satellites 1 and 3 are approximately 0.377 m apart. The desired relative positions are $d_{12}=0.42$ m, and $d_{13}=-0.45$ m.
Experiment 6.
This experiment demonstrates closed-loop repulsion between satellite 1 and 2, and closed-loop repulsion between satellite 1 and 3. Satellites 1 and 2 start at a relative distance of approximately 0.35 m, and satellites 1 and 3 start at a relative distance of approximately 0.38 m apart. The objective is to reach the desired relative positions $d_{12}=0.42$ m and $d_{13}=-0.45$ m. Figure 12 shows the experimental results, where the column on the left shows relative measurements, estimates, and average approximate force between satellite 2 (left satellite) and satellite 1 (middle satellite), as measured and computed on satellite 2; and the second column from the left shows the data as measured and computed on satellite 1. The third column from the left shows relative measurements, estimates, and average approximate force between satellite 1 and satellite 3 (right satellite), as measured and computed on satellite 1; and the column on the right shows the data as measured and computed on satellite 3. Figure 12 shows that $\hat{r}_{12}$ and $\hat{r}_{13}$ converges to $d_{12}$ and $d_{13}$ at approximately $t=30$ s.
Next, we present simulations of the closed-loop 3 and 4 where $\mathbf{p}_{ij,k}=p_{ij,k}\mathbf{i}$. The effect of friction between the satellites and the linear air tracks becomes noticeable with the addition of the third satellite. To model this effect, we include linear damping $\hat{B}_{i}v_{i}$ in 3, where $\hat{B}_{i}=0.08$ N–s/m. The simulated relative position includes zero-mean Gaussian-white sensor noise with variance $V=2\times 10^{-6}$ m2, which is the sensor noise variance estimated from experimental data. Figure 13 shows the simulation results, which qualitatively agree with the experimental results. However, we observe an oscillatory transient response in Figure 13, which is not observed in Figure 12. This can partly be explained by the imperfect friction model, unmodeled external disturbances (e.g., lateral air pressure from the air track), and unmodeled near-field electromagnetic effects. $\triangle$
Experiment 7.
This experiment demonstrates closed-loop attraction between satellite 1 and satellite 2, and closed-loop attraction between satellite 1 and satellite 3. Satellites 1 and 2 start at a relative distance of approximately 0.425 m, and satellites 1 and 3 start at a relative distance of approximately 0.46 m. The objective is to reach the desired relative positions $d_{12}=0.35$m and $d_{13}=-0.38$ m. Figure 14 shows that $\hat{r}_{12}$ and $\hat{r}_{13}$ converges to the desired relative positions $d_{12}$ and $d_{13}$ by approximately $t=26$ s, which demonstrates that the satellites achieve formation. Figure 15 shows the simulation results, which qualitatively agree with the experimental results. Similar to the comparison in 6, the simulation has oscillatory overshoot whereas the overshoot in the experiment decays more quickly. $\triangle$
Figure 14: Experiment demonstrating closed-loop attraction between satellite 1 and 2 and closed-loop attraction between satellite 1 and 3. Initially, satellites 1 and 2 are approximately 0.425 m apart, and satellites 1 and 3 are approximately 0.46 m apart. The desired relative positions are $d_{12}=0.35$ m, and $d_{13}=-0.38$ m.
Figure 15: Simulation demonstrating closed-loop attraction between satellite 1 and 2 and closed-loop attraction between satellite 1 and 3. Initially, satellites 1 and 2 are approximately 0.425 m apart, and satellites 1 and 3 are approximately 0.46 m apart. The desired relative positions are $d_{12}=0.35$ m, and $d_{13}=-0.38$ m.
Experiment 8.
This experiment demonstrates closed-loop attraction between satellite 1 and satellite 2, and closed-loop repulsion between satellite 1 and satellite 3. Satellites 1 and 2 start at a relative distance of approximately 0.46 m, and satellites 1 and 3 start at a relative distance of approximately 0.38 m apart.The objective is to reach the desired relative positions $d_{12}=0.42$ m and $d_{13}=-0.45$ m. Figure 16 shows that $\hat{r}_{12}$ and $\hat{r}_{13}$ converges to the desired relative position $d_{12}$ and $d_{13}$ by approximately $t=25$ s. Figure 17 shows the simulation of the experiment, which qualitatively matches the experimental results in Figure 16. $\triangle$
Figure 16: Experiment demonstrating closed-loop attraction between satellite 1 and 2 and closed-loop repulsion between satellite 1 and 3. Initially, satellites 1 and 2 are approximately 0.46 m apart, and satellites 1 and 3 are approximately 0.38 m apart. The desired relative positions are $d_{12}=0.42$m and $d_{13}=-0.45$ m.
Figure 17: Simulation demonstrating closed-loop attraction between satellite 1 and 2 and closed-loop repulsion between satellite 1 and 3. Initially, satellites 1 and 2 are approximately 0.46 m apart, and satellites 1 and 3 are approximately 0.38 m apart. The desired relative positions are $d_{12}=0.42$m and $d_{13}=-0.45$ m.
8 Concluding Remarks
This article presented a 3-satellite experimental demonstration of decentralized EMFF using AMFF. To our knowledge, this is the first demonstration of AMFF with at least 3 satellites in open or closed loop. The experimental results in this article not only demonstrated the feasibility of EMFF with AMFF but also demonstrated that the frequency-multiplexed amplitude-modulated sinusoidal approach can be used to achieve formation with multiple satellites. Future experiments could include 3 degree-of-freedom experiments (planar translation and attitude control) for 2 or more satellites.
Conflict of Interest Statement
There is no conflict of interest.
Acknowledgments
This work is supported in part by the National Aeronautics and Space Administration (80NSSC17M0040), the National Science Foundation (1932105), and the Air Force Office of Scientific Research (FA9550-20-1-0028).
We would like to thank Floyd Taylor, Herb Mefford, and Jonathan T. Williams at the University of Kentucky for their invaluable advice on improving the physical testbed.
Appendix A Proof of Proposition 2
Proof of Proposition 2. We consider 2 cases: (i) $\mathbf{r}\times\mathbf{f}_{*}\neq\mathbf{0}$ and (ii) $\mathbf{r}\times\mathbf{f}_{*}=\mathbf{0}$.
First, consider (i) $\mathbf{r}\times\mathbf{f}_{*}\neq\mathbf{0}$, and substituting $|\mathbf{r}\times\mathbf{f}_{*}|=|\mathbf{r}|^{2}\mathbf{f}_{*}-(\mathbf{r}\cdot\mathbf{f}_{*})\mathbf{r}$, (16), and (17) into (2) we get
$\mathbf{f}(\mathbf{r},\mathbf{g},\mathbf{h})=$ $\mathopen{}\mathclose{{\left(-\frac{2g_{\mathrm{r}}h_{\mathrm{r}}}{|\mathbf{r}|}+\frac{g_{\mathrm{rf}}h_{\mathrm{rf}}}{|\mathbf{r}|}-\frac{(g_{\mathrm{r}}h_{\mathrm{rf}}+g_{\mathrm{rf}}h_{\mathrm{r}})(\mathbf{r}\cdot\mathbf{f}_{*})}{|\mathbf{r}||\mathbf{r}\times\mathbf{f}_{*}|}}}\right)\mathbf{r}$
$+\frac{(g_{\mathrm{r}}h_{\mathrm{rf}}+g_{\mathrm{rf}}h_{\mathrm{r}})|\mathbf{r}|}{|\mathbf{r}\times\mathbf{f}_{*}|}\mathbf{f}_{*}.$ (40)
where the arguments $\mathbf{r}$ and $\mathbf{f}_{*}$ are omitted for brevity. We consider 2 subcases: $\mathbf{r}\cdot\mathbf{f}_{*}=0$ and $\mathbf{r}\cdot\mathbf{f}_{*}\neq 0$. For $\mathbf{r}\cdot\mathbf{f}_{*}=0$, it follows from (10)–(13) that $g_{\mathrm{r}}=h_{\mathrm{rf}}=0$ and $g_{\mathrm{rf}}h_{\mathrm{r}}=|\mathbf{r}\times\mathbf{f}_{*}|/|\mathbf{r}|$, and substituting into (40) yields $\mathbf{f}(\mathbf{r},\mathbf{g},\mathbf{h})=\mathbf{f}_{*}$. Next, for $\mathbf{r}\cdot\mathbf{f}_{*}\neq 0$, it follows from (14) and (15) that $\Phi_{2}=\Phi_{1}$, and using (10)–(13) yields
$g_{\mathrm{r}}h_{\mathrm{r}}$ $=-\frac{\operatorname{sgn}(\mathbf{r}\cdot\mathbf{f}_{*})}{4|\mathbf{r}|}\mathopen{}\mathclose{{\left(|\mathbf{r}\cdot\mathbf{f}_{*}|+\Phi_{1}}}\right),$
$g_{\mathrm{r}}h_{\mathrm{rf}}$ $=g_{\mathrm{rf}}h_{\mathrm{r}}=\frac{|\mathbf{r}\times\mathbf{f}_{*}|}{2|\mathbf{r}|}$
$a_{y}b_{y}$ $=-\frac{\operatorname{sgn}(\mathbf{r}\cdot\mathbf{f}_{*})}{2|\mathbf{r}|}\mathopen{}\mathclose{{\left(-|\mathbf{r}\cdot\mathbf{f}_{*}|+\Phi_{1}}}\right),$
Thus, substituting into (40) yields $\mathbf{f}(\mathbf{r},\mathbf{g},\mathbf{h})=\mathbf{f}_{*}$, which confirms the result for (i) $\mathbf{r}\times\mathbf{f}_{*}\neq 0$.
Next, we consider (ii) $\mathbf{r}\times\mathbf{f}_{*}=\mathbf{0}$, and substituting (10)– (13), (16), and (17) into (2) yields
$\mathbf{f}(\mathbf{r},\mathbf{g},\mathbf{h})$ $=-2g_{\mathrm{r}}h_{\mathrm{r}}\frac{\mathbf{r}}{|\mathbf{r}|}$
$=\frac{\operatorname{sgn}(\mathbf{r}\cdot\mathbf{f}_{*})}{2|\mathbf{r}|^{2}}(|\mathbf{r}\cdot\mathbf{f}_{*}|+\Phi_{1})^{\frac{1}{2}}(|\mathbf{r}\cdot\mathbf{f}_{*}|+\Phi_{2})^{\frac{1}{2}}\mathbf{r}.$ (41)
We consider 2 subcases: $\mathbf{r}\cdot\mathbf{f}_{*}=0$ and $\mathbf{r}\cdot\mathbf{f}_{*}\neq 0$. For $\mathbf{r}\cdot\mathbf{f}_{*}=0$, it follows from (41) that $\mathbf{f}(\mathbf{r},\mathbf{g},\mathbf{h})=\mathbf{0}$. Since $\mathbf{r}\cdot\mathbf{f}_{*}=0$ and $\mathbf{r}\times\mathbf{f}_{*}=\mathbf{0}$, it follows that $\mathbf{f}_{*}=\mathbf{0}$, which implies that $\mathbf{f}(\mathbf{r},\mathbf{g},\mathbf{h})=\mathbf{f}_{*}$. Next, for $\mathbf{r}\cdot\mathbf{f}_{*}\neq 0$, it follows from (14) and (15) that $\Phi_{2}=\Phi_{1}$. Since $\mathbf{r}\cdot\mathbf{f}_{*}\neq 0$ and $\mathbf{r}\times\mathbf{f}_{*}=\mathbf{0}$, it follows that $\mathbf{r}=(\operatorname{sgn}(\mathbf{r}\cdot\mathbf{f}_{*})|\mathbf{r}|/|\mathbf{f}_{*}|)\mathbf{f}_{*}$ and $|\mathbf{r}\cdot\mathbf{f}_{*}|+\Phi_{1}=2|\mathbf{r}||\mathbf{f}_{*}|$. Thus, substituting into (41) yields $\mathbf{f}(\mathbf{r},\mathbf{g},\mathbf{h})=\mathbf{f}_{*}$, which confirms the result for (ii) $\mathbf{r}\times\mathbf{f}_{*}=\mathbf{0}$. $\square$